# Trapped fireshell (halo) of photons and pairs around black-hole horizon: source for ultra-high-energy particles

### v1

@@ 3.3 Drifting electrons in an opaque fluid of photons and pairs
where the average $\langle\cdot\cdot\cdot\rangle_{a}$ is over the distribution of bulk and drifting electrons. The drifting electrons should be a small fraction of the bulk electrons. The same discussions apply to drifting $N_{p}$ and bulk $\bar{N}_{p}$ protons in the opaque fluid of photons and pairs.
We further examine the time scale $\tau_{a}^{e}$ of the electron acceleration process and compare it with the thermalisation time scale $\tau_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}c$ (2.7) of photon and pair fluid. Using Eq. (3.2) we define
@@
where the gravitational time dilation factors $d\tau=(1-1/r)^{1/2}dt$ are simplified. Recall that the thermalisation time scale $\tau_{\gamma}$ (2.7) represents also the time scale of the flux $F_{\gamma}$ of multi-photons acting on a changed particle.
@@
Moreover, due to the large opacity and frequent collisions with photons and pairs, drifting electrons ($\gamma_{e}\gtrsim 1$) quickly dissipate their energy and run back to bulk electrons ($\gamma_{e}=1$), which are in an energy equipartition with photons and pairs. It implies that the perturbations of drifting electrons ($\gamma_{e}>1$) should be rapidly washed out. Therefore, it looks unlikely that in the opaque fluid of photons and pairs, such small perturbations could develop, namely, accelerated electrons by the Compton-rocket effect could drift to a large Lorentz factor $\gamma_{e}\gg 1$, unless some instabilities of avalanche phenomena occur.
4 Runaway electrons in an opaque fluid photons and pairs
4.1 Accelerated electron runaway due to Klein–Nishina corrections
As discussed, in the opaque fluid of photons and pairs, the Compton-rocket effect on accelerating electrons is an approximately “local” and “instantaneous” perturbation that triggers electrons to increase energy and drift in the outward radial direction. On their way, accelerated electrons collide with photons and pairs with the probability described by the Klein–Nishina (KN) cross section (4.1), which, however, decreases as the accelerated electrons’ energy increases, and the straightforward scattering with a narrower deflection angle $\theta_{e}\sim 0$. This property of KN corrections implies the instability of an unstable avalanche runaway process. Namely, accelerated electrons gain more energy and have fewer collisions. Thus, they have greater chances to drift a longer distance and gain even more energy from the Compton-rocket effect. As a result, there is a probability that a small fraction of drifting electrons run away and achieve ultra-high energies, penetrating through the opaque fluid of photons and pairs.
@@ 4.1 Accelerated electron runaway due to Klein–Nishina corrections
where the factor $2$ in $\omega_{e}$ (4.1) is due to the blue-shift effect [30]. The cross-section decreases as the drifting electrons’ Lorentz factor $\gamma_{e}$ increases and vanishes when $\gamma_{e}\gg 1$. Therefore, high $\gamma_{e}$ drifting electrons have a small probability of being scattered back to bulk electrons with $\gamma_{e}=1$. Instead, they are further accelerated to higher $\gamma_{e}$ by the Compton-rocket effect, and behave as runaway electrons.
@@
and the gravitational time dilation factors $d\tau=(1-1/r)^{1/2}dt$ are simplified. The time scales (3.5), (3.6) and (4.2) give $\tau^{e}_{a}<\tau_{\gamma}\ll\tau^{e}_{{}_{\rm KN}}$, implying that rather than being smeared out, the instability could develop a runaway solution. It makes drifting electrons start from microscopic perturbations (3.6) run to macroscopic motions (4.2). It is necessary to perform numerical simulations to verify that such an instability and runaway solution indeed occur in an opaque fluid of photons and pairs at high temperature $T_{\gamma}\gg m_{e}$.
4.2 Electrons accelerated by photons and pairs at high temperature
Even if this instability occurs, what is the significant probability of runaway drifting electrons that survive over a long macroscopic distance and time scale, reaching ultra-high energies? We describe this probability by the fraction $N_{e}(\gamma_{e})/\bar{N}_{e}$ of drifting and bulk electrons. To study this fraction, we rewrite the Compton-rocket effect on the electron acceleration (3.2) in a small distance $dr=\beta_{e}d\tau=\beta_{e}(1-r_{+}/r)^{1/2}dt$ as
@@ 4.2 Electrons accelerated by photons and pairs at high temperature
In the one-dimensional model, this equation represents drifting electrons $N_{e}(\gamma_{e})$ of $\gamma_{a}>1$ accelerated out from bulk electrons $\bar{N}_{e}$ of $\gamma_{e}=\langle\tilde{\gamma}_{e}\rangle=1$. The effective mass $m^{\rm eff}_{e}\approx m_{n}+m_{e}$ will be explained in due course.
First, we consider an unrealistic situation where all bulk electrons $\bar{N}_{e}$ (3.3) are accelerated to drifting electrons $N_{e}(\gamma_{e})$ in the radial direction without any collision with photons and pairs. Then integrating the Compton-rocket effect (4.3) yields
@@
where bulk electrons $\gamma_{e}=1$ near the horizon $r_{+}$ are accelerated to drifting electrons $\gamma_{e}>1$ at the radius $r$. The photon and pair flux $F_{\gamma}(r)=\rho_{\gamma}c$ is highly peaked and has a great gradient near the horizon $r_{+}$, as shown in Fig. 2. Therefore, the photon and pair flux $F_{\gamma}(r_{+})=\rho_{\gamma}(r_{+})c$ and temperature $T_{\gamma}(r_{+})$ (2.9) near the horizon predominantly contribute to the integration (4.4), i.e., to the acceleration of drifting electrons out from bulk electrons. Similarly to Eq. (2.5) for the total energy, we use the $\delta$-function approximation $\int(1-r_{+}/r)^{-1/2}\rho_{\gamma}(r)dr\approx\rho_{\gamma}(r_{+})\delta\ell=\sigma_{{}_{\rm SB}}d_{f}T^{4}_{\gamma}(r_{+})\delta\ell$, where $\delta\ell=(1-r_{+}/r)^{-1/2}dr$ is a small proper distance near the horizon. Solution (4.4) approximately becomes
@@
namely, the radiation force $F_{\gamma}(r_{+})=\rho_{\gamma}(r_{+})c$ accelerates on an electron over a distance $\delta\ell$ to reach its Lorentz factor $\gamma_{e}$.
@@
In the viewpoint of the energy density $\rho_{\gamma}\sim T_{\gamma}^{4}\sim E^{2}_{\rm eff}/(8\pi)$, the photons and pairs radiation force (pressure) $F_{\gamma}=c\rho_{\gamma}$ for the temperature $T_{\gamma}^{+}\equiv T_{\gamma}(r_{+})\sim{\mathcal{O}}(10^{1-2})$MeV is equivalent to an effective electric force $eE_{\rm eff}$, which is about ${\mathcal{O}}(10^{1-2})$ times larger than the critical electric field force $eE_{c}=m_{e}^{2}$ for electron-positron pair productions. In a classical scenario for a test charged particle accelerated by an electric field, we estimated that such an effective electric force $eE_{\rm eff}$ accelerates an electron to PeV $\sim 10^{9}m_{e}(\gamma_{e}\sim 10^{9})$ energies in a small distance $d\sim{\mathcal{O}}(10^{9})\lambda_{e}\sim{\mathcal{O}}(10^{-2})$ cm and time $\tau=d/c\sim{\mathcal{O}}(10^{-12})$ sec [24]. Therefore, the macroscopic length $\delta\ell$ determined from the total thermal energy $E_{\gamma}$ (2.10) should be large enough for an electron reaching $\gamma_{e}\sim 10^{9}$ in Eq. (4.5). It indicates that the macroscopic sizes and geometries of dense and energetic photon and pair spots are irrelevant for the Compton-rocket effect, because the local inhomogeneous density of photons and pairs accelerates electrons on microscopic time and length scales.
@@
Physically, the solution (4.5) without any collision means that (i) all bulk electrons $\bar{N}_{e}$ of $\gamma_{e}=1$ accelerated to drifting electrons $N_{e}(\gamma_{e})$ of an enormous Lorentz factor $\gamma_{e}\gg 1$; (ii) the majority of drifting electrons get accelerated and reach their ultra-high energies in the neighbourhood near the horizon. This collisionless solution (4.5) of $N_{e}(\gamma_{e})=\bar{N}_{e}$ is certainly not realistic, since the fluid of photons and pairs is extremely opaque.
4.3 Superthermal spectrum of UHE electrons
4.3.1 Thomason scattering in low energies
@@
Second, we consider the case of the constant Thomason cross section $\sigma_{{}_{T}}$ in low energies (temperatures). We calculate the number of drifting electrons accelerated in one proper mean-free path $\xi_{e}^{\rm T}=(n_{\gamma}\sigma_{{}_{T}})^{-1}$ and time $d\tau=\xi_{e}^{\rm T}/\beta_{e}$. The acceleration equation (3.2) for the Lorentz factor variation $\Delta\gamma_{e}$ in one mean-free path becomes
@@
Due to the Thomson scatterings with photons and pairs, the drifting electron numbers $N_{e}(\gamma_{e})$ decrease following the one-dimensional rate equation,
@@
$\frac{N_{e}(\gamma_{e})}{\bar{N}_{e}}$ $=$ $\exp-\int_{1}^{\gamma_{e}}\frac{1}{\Delta\gamma^{\rm T}_{e}}{d\gamma_{e}}\approx\left(\frac{3}{2\gamma_{e}^{2}+1}\right)^{2.78(m^{\rm eff}_{e}/T^{+}_{\gamma})}$. (4.10)
@@
Thus, the most accelerated drifting electrons $N_{e}(\gamma_{e})$ at high energies run back to the bulk electrons $\bar{N}_{e}$ at low energies, which are in the energy equipartition with photons and pairs. In other words, the probability of drifting electrons reaching a large Lorentz factor $\gamma_{e}\gg 1$ goes to zero. Very few electrons are in the high-energy superthermal spectrum tail. It is the case for the Sun, whose core temperature of photons is much smaller than the electron mass, i.e., $T_{\odot}\approx 1.5\times 10^{4}{\rm eV}\ll m_{e}$.
Figure 4: We calculate the fraction $N_{e}/\bar{N}_{e}$ of accelerated electrons drifting out of the bulk electrons for the Thomson (4.10) and the Klein-Nishina (4.15) scattering cases. It shows that the probability of accelerated electrons reaching high energies is non-trivial, implying the runaway behaviour, in the Klein-Nishina scattering case, but is completely negligible in the Thomson scattering case. These figures also present the proton case with the substitution of the subscript $e\rightarrow p$.
4.3.2 Klein-Nishina scattering in high energies
Finally, we consider drifting electrons accelerated in one proper mean-free path $\xi_{e}^{\rm KN}=(n_{\gamma}\sigma^{e}_{{}_{\rm KN}})^{-1}$, and show they run away due to the energy-dependent Klein-Nishina cross section $\sigma^{e}_{{}_{\rm KN}}(\gamma_{e})$ (4.1). The acceleration equation (3.2) for the Lorentz factor variation $\Delta\gamma^{\rm KN}_{e}$ in one mean-free path becomes

### v2

@@ 3.3 Drifting electrons versus bulk electrons
where the average $\langle\cdot\cdot\cdot\rangle_{a}$ is over the energy-momentum distribution of bulk and drifting electrons. The drifting electrons $N_{e}(\gamma_{e})$ should be a small fraction of the bulk electrons $\bar{N}_{e}$, because most of the radiation force accelerated electrons do not become drifting electrons, but run back to bulk electrons due to collisions with photons and pairs. We define the fraction (radio) of drifting and bulk electrons as
$\frac{N_{e}}{\bar{N}_{e}}\equiv\frac{N_{e}(\gamma_{e})}{\bar{N}_{e}},$ (3.4)
which actually represents the probability that electrons are accelerated out of bulk electrons to become drifting electrons by the Compton-rocket effect in an opaque photon-pair fluid. This fraction (3.4) is a crucial quantity we adopt to describe the significance of the Compton-rocket effect on accelerating electrons to a Lorentz factor $\gamma_{e}$ in an opaque photon-pair fluid. The same discussions apply to drifting $N_{p}$ and bulk $\bar{N}_{p}$ protons in the opaque fluid of photons and pairs.
We further examine the time scale $\tau_{a}^{e}$ of the electron acceleration process (Compton-rocket effect) and compare it with the thermalisation time scale $\tau_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}c$ (2.8) of photon and pair fluid. Using Eq. (3.2) we define
@@
Here, the gravitational time dilation factors $d\tau=(1-1/r)^{1/2}dt$ are simplified. Recall that the thermalisation time scale $\tau_{\gamma}$ (2.8) represents also the time scale of the radiation flux $F_{\gamma}$ of multi-photons acting on a changed particle.
@@
Moreover, due to the large opacity and frequent collisions with photons and pairs, drifting electrons ($\gamma_{e}\gtrsim 1$) quickly dissipate their energy and run back to bulk electrons ($\gamma_{e}=1$), which are in an energy equipartition with photons and pairs. It implies that the perturbations of drifting electrons ($\gamma_{e}>1$) should be rapidly washed out. Therefore, it looks unlikely that in the opaque fluid of photons and pairs, such small perturbations could develop, namely, some accelerated electrons by the Compton-rocket effect could drift to a large Lorentz factor $\gamma_{e}\gg 1$, unless some instabilities of avalanche phenomena occur.
4 Electrons runaway in an opaque radiation fluid
4.1 Instability and electron runaway due to Klein–Nishina corrections
As discussed, in the opaque fluid of photons and pairs, the Compton-rocket effect on accelerating electrons is an approximately “local” and “instantaneous” perturbation that triggers electrons to increase energy and drift in the outward radial direction. On their way, accelerated electrons collide with photons and pairs with the probability described by the Klein–Nishina (KN) cross section (4.1), which, however, decreases as the accelerated electrons’ energy increases, and the straightforward scattering with a narrower deflection angle $\theta_{e}\sim 0$. This property of KN corrections implies the instability of an unstable avalanche runaway process of drifting electrons. Namely, accelerated electrons gain more energy and have fewer collisions. Thus, they have greater chances to drift a longer distance and gain even more energy from the Compton-rocket effect. As a result, there is a probability that a small fraction of drifting electrons run away and achieve ultra-high energies, penetrating through the opaque fluid of photons and pairs.
@@ 4.1 Instability and electron runaway due to Klein–Nishina corrections
where the factor $2$ in $\omega_{e}$ (4.1) is due to the blue-shift effect [32]. The cross-section decreases as the drifting electrons’ Lorentz factor $\gamma_{e}$ increases and vanishes when $\gamma_{e}\gg 1$. Therefore, high $\gamma_{e}$ drifting electrons have a small probability of being scattered back to bulk electrons with $\gamma_{e}=1$. Instead, they are further accelerated to higher $\gamma_{e}$ by the Compton-rocket effect, and behave as runaway drifting electrons.
@@
and the gravitational time dilation factors $d\tau=(1-1/r)^{1/2}dt$ are simplified. The time scales (3.5), (3.6) and (4.2) give $\tau^{e}_{a}<\tau_{\gamma}\ll\tau^{e}_{{}_{\rm KN}}$, implying that rather than being smeared out, the instability could develop a runaway solution. It makes drifting electrons start from microscopic perturbations (3.6) and run to macroscopic motions (4.2).
4.2 Probability of drifting electrons runaway
Even if such an instability occurs, in view of drifting electrons colliding with photons and pairs, what is the probability that runaway drifting electrons survive when they travel over a long macroscopic distance and time scale, reaching a large Lorentz factor $\gamma_{e}$? We describe this probability by the fraction $N_{e}/\bar{N}_{e}$ (3.4) of drifting and bulk electron numbers, and study it for the following three cases:
  1. (i)
    $N_{e}/\bar{N}_{e}=1$ (maximal probability) corresponds to all of the bulk electrons being accelerated and becoming drifting electrons at Lorentz factor $\gamma_{e}\gg 1$. This collisionless case can not be realistic, since the drifting electrons dissipate their kinetic energies, and return to bulk electrons via collisions with photons and pairs in the fireshell of a large opacity. Nevertheless, we study this case to gain an insight into the nature of the Compton-rocket effect on accelerating electrons inside an opaque fluid of energetic and dense photons and pairs.
  2. (ii)
    $N_{e}/\bar{N}_{e}\approx 0$ (near zero probability) corresponds to almost none of the bulk electrons being accelerated and becoming drifting electrons at Lorentz factor $\gamma_{e}\gg 1$. In other words, almost all drifting electrons return to bulk electrons due to collisions with photons and pairs. It indicates that the instability does not develop into an avalanche runaway process, and a negligible amount of bulk electrons are accelerated and become drifting electrons at Lorentz factor $\gamma_{e}\gg 1$. It is the collisional case of Thomson scattering with an energy-independent cross section at low energies.
  3. (iii)
    $N_{e}/\bar{N}_{e}\not=0$ is a non-trivial value (finite probabilities), corresponding to a significant amount of the bulk electrons being accelerated and becoming drifting electrons at Lorentz factor $\gamma_{e}\gg 1$. It indicates that the instability has developed into an avalanche runaway process, and a significant amount of bulk electrons are accelerated and become drifting electrons at Lorentz factor $\gamma_{e}\gg 1$. It is the collisional case of Klein–Nishina scattering with an energy-dependent cross section at high energies.
To quantitatively study this fraction $N_{e}/\bar{N}_{e}$ in these cases, we need two master equations for drifting electrons: (1) the acceleration equation (4.3) attributed to the Compton-rocket effect, (2) the decay rate equation (4.8) or (4.13) due to drifting electrons’ energy-momentum dissipation in collisions with photons and pairs inside the fireshell.
We rewrite the Compton-rocket effect on the electron acceleration (3.2) in a small distance $dr=\beta_{e}d\tau=\beta_{e}(1-r_{+}/r)^{1/2}dt$ as
@@ 4.2 Probability of drifting electrons runaway
In the one-dimensional model, this equation represents drifting electrons $N_{e}(\gamma_{e})$ of $\gamma_{e}>1$ accelerated out of bulk electrons $\bar{N}_{e}$ of $\gamma_{e}=\langle\tilde{\gamma}_{e}\rangle=1$. The boundary condition is $\gamma_{e}=1$ at $r=r_{+}$ from the initial condition for (3.2). The effective mass $m^{\rm eff}_{e}\approx m_{p}+m_{e}$ will be explained in due course.
4.3 Electrons accelerated by an opaque radiation fluid at high temperature
To have an idea about the spacetime length scales of electron acceleration in the fireshell of energetic and dense photons and pairs, we first consider the collisionless case where all bulk electrons $\bar{N}_{e}$ (2.13) are accelerated to drifting electrons $N_{e}(\gamma_{e})$ in the radial direction without any collision with photons and pairs. Then integrating the Compton-rocket effect (4.3) yields
@@ 4.3 Electrons accelerated by an opaque radiation fluid at high temperature
where bulk electrons $\gamma_{e}=1$ near the horizon $r_{+}$ are accelerated to drifting electrons $\gamma_{e}>1$ at the radius $r$. The photon and pair flux $F_{\gamma}(r)=\rho_{\gamma}c$ is highly peaked and has a great gradient near the horizon $r_{+}$, as shown in Fig. 2. Therefore, the photon and pair flux $F^{+}_{\gamma}=\rho_{\gamma}(r_{+})c$ (2.12) and temperature $T^{+}_{\gamma}=T_{\gamma}(r_{+})$ (2.10) near the horizon predominantly contribute to the integration (4.4), i.e., to the acceleration of drifting electrons out of bulk electrons. Similarly to Eq. (2.6) for the total energy, we use the $\delta$-function approximation $\int(1-r_{+}/r)^{-1/2}\rho_{\gamma}(r)dr\approx\rho_{\gamma}(r_{+})\delta\ell=\sigma_{{}_{\rm SB}}d_{f}T^{4}_{\gamma}(r_{+})\delta\ell$, where $\delta\ell=(1-r_{+}/r)^{-1/2}dr$ is a small proper distance near the horizon. Solution (4.4) approximately becomes
@@
namely, the radiation force $F_{\gamma}(r_{+})=\rho_{\gamma}(r_{+})c$ near the horizon, where the energy density $\rho_{\gamma}(r_{+})$ and temperature $T^{+}_{\gamma}$ are maximal, accelerates on an electron over a distance $\delta\ell$ to reach its Lorentz factor $\gamma_{e}$.
@@
In the viewpoint of the energy density $\rho_{\gamma}\sim T_{\gamma}^{4}\sim E^{2}_{\rm eff}/(8\pi)$, the photons and pairs radiation force (pressure) $F_{\gamma}=c\rho_{\gamma}$ for the temperature $T_{\gamma}^{+}\equiv T_{\gamma}(r_{+})\sim{\mathcal{O}}(10^{1-2})$MeV is equivalent to an effective electric force $eE_{\rm eff}$, which is about ${\mathcal{O}}(10^{1-2})$ times larger than the critical electric field force $eE_{c}=m_{e}^{2}$ for electron-positron pair productions. In a classical scenario of a test charged particle accelerated by an electric field, we estimated that such an effective electric force $eE_{\rm eff}$ accelerates an electron to PeV $\sim 10^{9}m_{e}(\gamma_{e}\sim 10^{9})$ energies in a small distance $d\sim{\mathcal{O}}(10^{9})\lambda_{e}\sim{\mathcal{O}}(10^{-2})$ cm and time $\tau=d/c\sim{\mathcal{O}}(10^{-12})$ sec [17]. Therefore, the macroscopic length $\delta\ell$ determined from the total thermal energy $E_{\gamma}$ (2.11) should be large enough for an electron reaching $\gamma_{e}\sim 10^{9}$ in Eq. (4.5). It indicates that the macroscopic sizes and geometries of dense and energetic photon-pair fireshells or spots are irrelevant for the Compton-rocket effect, because the local inhomogeneous density of photons and pairs accelerates electrons on microscopic time and length scales $d$, which are much smaller than the macroscopic time and size variation of fireshells or firespots.
@@
Physically, the collisionless solution (4.5) means that (i) all bulk electrons $\bar{N}_{e}$ of $\gamma_{e}=1$ accelerated to drifting electrons $N_{e}(\gamma_{e})$ of an enormous Lorentz factor $\gamma_{e}\gg 1$; (ii) the majority of drifting electrons get accelerated and reach their ultra-high energies in the neighbourhood near the horizon, or where photon-pair energy density and temperature are maximal. This collisionless solution (4.5) of $N_{e}(\gamma_{e})=\bar{N}_{e}$ is certainly not realistic, since the radiation fluid of photons and pairs is extremely opaque, and drifting electrons colliding with photons and pairs is unavailable. However, the study of the collisionless case (4.5) let us learn (i) the large-temperature $T^{+}_{\gamma}$ dominates the radiation-force contribution to electron acceleration; (ii) the spacetime length scales $d$ of electron acceleration in an opaque radiation fluid of photons and pairs. This helps make appropriate approximations in studying collisional cases, in which the drifting electron cannot reach $\gamma_{e}\gg 1$ (4.5) with the maximal probability (fraction) $N_{e}/\bar{N}_{e}=1$.
For the collisional case, to show the runaway dynamics due to the Klein-Nishina correction in high energies, as previously speculated in section (4.1), we attempt to calculate the fraction $N_{e}/\bar{N}_{e}$ numerically in two separate cases: (1) the stable case, where drifting electrons undergo Thomson scattering in low energies; (2) the unstable case, where drifting electrons undergo Klein-Nishina scattering in high energies. We do not adopt the full scattering cross-section that varies with energy. The reasons are that (i) there is no analytical solution $N_{e}/\bar{N}_{e}$ for the entire energy range; (ii) we have numerical difficulties to tackle the runway dynamics when it occurs, the fraction $N_{e}/\bar{N}_{e}$ varies by many orders of magnitude, as will be shown in Figs. 4 and 5; (iii) it is helpful to distinctly compare and contrast Thomson and Klein-Nishina scattering cases to understand how the runaway dynamics physically develops.
4.4 Stable case: Thomson scattering at low energies
@@
Second, we consider the collisional case of the constant Thomason scattering cross section $\sigma_{{}_{T}}$ at low energies (temperatures). We calculate the number of drifting electrons accelerated in one proper mean-free path $\xi_{e}^{\rm T}=(n_{\gamma}\sigma_{{}_{T}})^{-1}$ and time $d\tau=\xi_{e}^{\rm T}/\beta_{e}$. The acceleration equation (3.2) or (4.3) for the Lorentz factor variation $\Delta\gamma^{\rm T}_{e}$ in one mean-free path becomes
@@
Due to the Thomson scatterings with photons and pairs, the drifting electron number $N_{e}(\gamma_{e})$ decreases following the one-dimensional rate equation,
@@
$\frac{N_{e}(\gamma_{e})}{\bar{N}_{e}}$ $=$ $\exp\left(-\int_{1}^{\gamma_{e}}\frac{1}{\Delta\gamma^{\rm T}_{e}}{d\gamma_{e}}\right)\approx\left(\frac{3}{2\gamma_{e}^{2}+1}\right)^{2.78(m^{\rm eff}_{e}/T^{+}_{\gamma})}$. (4.10)
@@
Thus, the most accelerated drifting electrons $N_{e}(\gamma_{e})$ at high energies run back to the bulk electrons $\bar{N}_{e}$ at low energies, which are in the energy equipartition with photons and pairs. In other words, the probability of drifting electrons reaching a large Lorentz factor $\gamma_{e}\gg 1$ goes to zero at low temperature $T^{+}_{\gamma}\ll m^{\rm eff}_{e}$. Very few electrons are in the high-energy superthermal spectrum tail; thus, they are negligible and irrelevant for observations of UHE particles. It is the case for the Sun, whose core temperature of photons is much smaller than the electron mass, i.e., $T_{\odot}\approx 1.5\times 10^{4}{\rm eV}\ll m_{e}$ and the fraction (4.10) is exponentially suppressed.
Figure 4: We calculate the fraction (probability) $N_{e}/\bar{N}_{e}$ of accelerated electrons drifting out of the bulk electrons for the Thomson (4.10) and the Klein-Nishina (4.15) scattering cases. It shows that the probability of accelerated electrons reaching high energies is non-trivial, implying the runaway behaviour, in the Klein-Nishina scattering case, but is completely negligible in the Thomson scattering case. These figures also represent the proton case with the substitution of the subscript $e\rightarrow p$.
4.5 Unstable case: Klein-Nishina scattering at high energies
Third, we investigate the collisional case of Klein-Nishina scattering at high energies (temperatures), and show a non-trivial probability of drifting electrons running away, reaching $\gamma_{e}\gg 1$, due to the energy-dependent nature of the KN cross section $\sigma^{e}_{{}_{\rm KN}}(\gamma_{e})$ (4.1), which decreases with increasing energies.
Analogously to $\Delta\gamma^{\rm T}_{e}$ (4.6), define $\Delta\gamma^{\rm KN}_{e}$ to be the Lorentz factor variation of drifting electrons accelerated in one proper mean-free path $\xi_{e}^{\rm KN}=(n_{\gamma}\sigma^{e}_{{}_{\rm KN}})^{-1}$. The acceleration equation (3.2) or (4.3) for the Lorentz factor variation $\Delta\gamma^{\rm KN}_{e}$ in one mean-free path $\xi_{e}^{\rm KN}$ becomes